# When Geoscience Meets Foundation Models: Towards General Geoscience Artificial Intelligence System

### v3

@@ III-A2 Adaptation
Datasets[85] Image-text datasets - - Remote sensing RSICap and RSIEval datasets for VLMs
III-B Large Language Models
Large language models (LLMs) have been widely used in many applications since BERT [86] was introduced. In the geoscience domain, LLMs have been used for various tasks, including geoscience question answering, query-based summarization, keyword classification, and geoscientific text data analysis.
Several LLMs for geoscience have been developed based on BERT [86], SciBERT [58], and LLaMA [87]. For example, Denli et al. [56] retrained a pre-trained BERT model with our 20M internal geoscientific records, referring to GeoBERT. They fine-tuned the GeoBERT model for several tasks including geoscience question answering and query-based summarization. The work [57] introduced BERT-E, an Earth science-specific language model built using transfer learning from SciBERT [58], which shows improved performance in keyword classification and highlights the need for robust benchmarks in evaluating such domain-specific language models, with potential applications across various scientific fields. The authors in [21] introduced K2, the first LLM specifically for geoscience, along with a new instruction tuning dataset, GeoSignal, and a benchmark, GeoBench, to adapt and evaluate LLMs in geoscience, demonstrating its effectiveness in understanding and utilizing geoscience knowledge with open-source data and model checkpoints.
Despite general geoscience tasks, LLMs have been used in specific geoscience domains. For example, the authors in [61] presented OceanGPT, the first-ever Large Language Model specialized for ocean science, developed with a novel framework, DoInstruct, for generating domain-specific instruction data and evaluated using OceanBench, a new oceanography benchmark, demonstrating enhanced expertise and preliminary embodied intelligence capabilities in ocean science tasks.
Additionally, the authors in [62] described the development and validation of new geoscience-specific language models trained on public geoscientific reports and publications, which outperform generic models in various NLP tasks and demonstrate the potential of using NLP to extract meaningful semantic relationships and latent knowledge from the vast volume of geoscience literature. [63] described the application of natural language processing (NLP) to geoscientific text data to extract meaningful semantic relationships and predict mineral potential, demonstrating how geoscience language models can transform unstructured text into actionable insights to streamline mineral exploration for critical materials. [64] presented the development of GeoGalactica, the largest language model specialized in geoscience, which is trained on a vast geoscience text corpus and fine-tuned with instruction-tuning data, highlighting the model’s comprehensive training process and its open-source contribution to the field.
In summary, the development of LLMs for geoscience has been an active area of research, with several models developed and evaluated for various geoscience tasks. These models have demonstrated the potential of using LLMs to extract meaningful insights from geoscience text data and to improve the performance of geoscience NLP tasks.
III-C Large Vision Models
Large vision models (LVMs) have been widely used in remote sensing and geoscience for various tasks, including remote sensing image analysis, satellite image classification, and climate and weather modeling. Most of the LVMs are based on the transformer architecture, which is effective in handling large-scale remote sensing data. For example, the authors in [14] introduced RingMo, a remote sensing (RS) foundation model framework using generative self-supervised learning, trained on a large-scale dataset of two million RS images to address the domain gap and generalization issues in RS image interpretation, showing state-of-the-art performance across multiple datasets and tasks, advocating for the adoption of generative self-supervised learning in RS research. The work [15] introduced LVMs specifically tailored for remote sensing tasks, featuring a new rotated varied-size window attention mechanism in vision transformers, demonstrating superior performance in detection, classification, and segmentation tasks with improved computational efficiency and data transferability. Cha et al. [65] discussed the impact of increasing the number of model parameters on the performance of foundation models in remote sensing tasks like rotated object detection and semantic segmentation, demonstrating that larger models improve performance and data efficiency, achieving state-of-the-art results on several benchmark datasets. The authors in [66] presented RingMo-Sense, a remote sensing foundation model for spatiotemporal prediction, utilizing spatial, temporal, and spatiotemporal triple-branch prediction networks with parameter-sharing and progressive joint training strategies, validated on an RS spatiotemporal dataset and demonstrating competitive performance across six downstream tasks. [67] introduced RingMo-lite, a lightweight CNN-Transformer hybrid network optimized for RS image interpretation on edge devices, effectively handling high- and low-frequency spectral components and reducing computational resources while maintaining high accuracy and state-of-the-art performance in various RS tasks, soon to be integrated into the MindSpore computing platform.
Moreover, several LVMS for geoscience are based on the Segment Anything Model (SAM), which is a transformer-based model that is effective in handling remote sensing data. For example, Yan et al. [52] introduced RingMo-SAM, a foundation model for multimodal remote-sensing image segmentation, capable of segmenting and identifying object categories in both optical and SAR data, using a large-scale dataset and a novel category-decoupling mask decoder and prompt encoder, demonstrating effectiveness across various datasets. The authors in [68] introduced SAMRS, a large-scale remote sensing segmentation dataset created using the Segment Anything Model (SAM) and existing RS object detection datasets, containing over 105,090 images and 1.6 million instances, with a comprehensive analysis and preliminary experiments demonstrating its utility for various segmentation and detection tasks, addressing challenges of limited training data. [36] presented the development of RSPrompter, an automated instance segmentation approach for remote sensing images based on the Segment Anything Model (SAM), which leverages prompt learning to improve SAM’s ability to generate semantically meaningful segmentation results, validated by extensive experiments on various datasets, with the code publicly available. Ma et al. [69] described a novel framework utilizing the Segment Anything Model (SAM) for semantic segmentation in remote sensing, introducing SAM-Generated Object (SGO) and SAM-Generated Boundary (SGB) concepts along with novel object and boundary loss components, demonstrating improved performance on two datasets, with the source code available online. The authors in [70] introduced Geographical SAM (GeoSAM), an advanced SAM-based framework enhanced with dense and sparse visual prompts from zero-shot learning and a pre-trained CNN, which significantly outperforms existing methods in segmenting mobility infrastructure in geographical images, marking a significant advancement in the use of foundation models for this purpose. The work [71] exploited OFA-Net (One-For-All Network), a versatile method using a single Transformer backbone for multiple data modalities and spatial resolutions in remote sensing, pre-trained on a multi-modal dataset, showing promising performance across 12 different downstream tasks, moving towards a unified foundation model in Earth vision.
For satellite image classification, the authors in [72] introduced CSP, a contrastive pre-trained model for satellite image classification, which outperforms existing methods in satellite image classification tasks, demonstrating the effectiveness of contrastive pre-training in remote sensing tasks. Similarly, [16] introduced Prithvi, a transformer-based geospatial foundational model pre-trained on over 1TB of multispectral satellite imagery, demonstrating its effectiveness in various Earth observation tasks through efficient pre-training and fine-tuning, with significant improvements over existing models and the ability to maintain accuracy with limited labeled data, made publicly available as an open-source contribution.
The authors in [73] presented ClimaX, a flexible and generalizable deep learning model for weather and climate science, extending the Transformer architecture to handle heterogeneous datasets and various spatio-temporal scales, demonstrating superior performance in weather forecasting and climate projections compared to other data-driven models, with the source code available online. The work [74] introduced W-MAE, a novel weather and climate model based on the Transformer architecture, which outperforms existing models in weather forecasting and climate change prediction, demonstrating the potential of large vision models in weather and climate science. Similarly, Fourcastnet [9], GraphCast [10], Fengwu [11], and Pangu-Weather [8] are all large vision models specifically tailored for weather forecasting and climate modeling, demonstrating the potential of large vision models in weather and climate science. Guo et al. citeguo2023skysense exploited SkySense, a large vision model for Earth image analysis, which outperforms existing models in Earth image analysis tasks, demonstrating the potential of large vision models in Earth image analysis.
To summarize, LVMs have been widely used in remote sensing and geoscience for various tasks, including remote sensing image analysis, satellite image classification, and climate and weather modeling. These models have demonstrated the potential of using large vision models to improve the performance of remote sensing and geoscience tasks.
III-D Large Vision-Language Models
Large vision-language models (LVLMs) are increasingly being used in the fields of remote sensing and geoscience. LVLMs are often utilized for visual question answering. The authors in [76] first described how Remote Sensing Visual Question Answering (RSVQA) leverages advancements in natural language processing and computer vision to analyze satellite images with language processing, comparing the effectiveness of different language models, particularly focusing on the benefits of fine-tuning in enhancing performance. The work [77] presented LiT-4-RSVQA, a lightweight transformer-based architecture for efficient and accurate visual question answering in remote sensing, which reduces computational demands while maintaining accuracy, as demonstrated by its performance on a VQA benchmark dataset. Chappuis et al. [78] outlined how Remote Sensing Visual Question Answering (RSVQA) facilitates human-machine interaction with overhead imagery using natural language, emphasizing the need to address language biases in RSVQA through a comprehensive analysis involving visual blind models, adversarial testing, and dataset analysis, revealing more severe biases in remote sensing compared to standard VQA due to dataset specifics.
Additionally, LVLMs have been used for remote sensing image understanding. Kuckreja et al. [79] introduced GeoChat, the first remote sensing Vision-Language Model with multitask conversational capabilities, designed to overcome the limitations of general-domain VLMs in handling high-resolution, category-diverse remote sensing imagery. GeoChat excels in a variety of RS tasks and addresses the challenge of a lack of domain-specific datasets by creating a novel RS multimodal instruction-following dataset. The authors in [80] outlined EarthGPT, a large vision-language model for Earth observation data, which outperforms existing models in Earth observation tasks, demonstrating the potential of large vision-language models in Earth observation. Zhan et al. [81] introduced SkyeyeGPT, a large vision-language model for remote sensing, which outperforms existing models in remote sensing tasks, demonstrating the potential of large vision-language models in remote sensing. The work [85] discussed the development of a high-quality Remote Sensing Image Captioning dataset (RSICap) and a benchmark evaluation dataset (RSIEval) to advance VLMs in remote sensing, addressing the lack of comprehensive image-text datasets for training and effectively evaluating these models in the domain. Mall et al. [82] introduced a method training LVLMs for remote-sensing images without textual annotations by aligning an image encoder for remote sensing images with CLIP’s image encoder using paired internet and satellite images, enabling zero-shot, open-vocabulary image classification, retrieval, segmentation, and visual question answering, outperforming supervised models on several tasks. Similarly, Li et al. [83] introduced RS-CLIP, an LVLM for remote sensing scene classification that uses contrastive vision-language supervision and pseudo-labeling techniques, achieving high zero-shot classification accuracy on four benchmark datasets by learning semantic-aware visual representations and applying a curriculum learning strategy.
Zhang et al. [84] discussed the evaluation of LVLMs on Earth observation data through a comprehensive benchmark focusing on scene understanding, localization, counting, and change detection tasks, revealing that while these models excel in tasks like location understanding and captioning, their spatial reasoning limitations affect performance in object localization and counting. The authors in [85] discussed the development of a high-quality Remote Sensing Image Captioning dataset (RSICap) and a benchmark evaluation dataset (RSIEval) to advance LVLMs in remote sensing, addressing the lack of comprehensive image-text datasets for training and effectively evaluating these models in the domain.
LVLMs have been widely used in remote sensing and geoscience for various tasks, including visual question answering and remote sensing image understanding. These models have demonstrated the potential of using large vision-language models to improve the performance of remote sensing and geoscience tasks.
IV Challenges of foundation models in Geoscience AI
IV-A Validation and Verification
GFMs possess the unique capacity to process complex inputs and generate likewise complex outputs. This makes validating their correctness exceptionally challenging, if not impossible. Unlike conventional geoscience AI models designed for specific tasks with supervised learning approaches validated within predefined use case parameters, GFMs promise to perform novel tasks specified by end-users without further training. Therefore, anticipating all possible failure modes for such broadly capable systems a prior becomes categorically more difficult. Consequently, developers and regulators bear the responsibility of transparently delineating how GFMs have been tested and which use cases they have been approved to operate within. Comprehensive testing methodologies and use case definitions will be critical for ensuring proper functioning and, hence, trust.
Many geoscience processes operate over prolonged timescales, including climate change [8, 7], ice sheet dynamics, or tectonic motions, necessitating consistent Earth observations spanning decades to centuries or even-in that case from proxy data at paleoclimate time scales, to sufficiently characterize them. Testing GFMs requires amassing illustrative cases over comparably long time periods to encapsulate rare events statistically. Even within a human lifespan, acquiring robust statistical samples poses challenges. Furthermore, when foundation models fail at predicting major events, responding systematically to identify specific weaknesses proves more constructive than overreacting to isolated failures. Often, targeted model refinements emerge from rigorous post-event analyses, contingent on collecting appropriate data preceding the event. Adopting a measured, evidence-based approach for incrementally improving models through their occasional failures will promote greater confidence and utility over the long term.
IV-B Multi-Source Data
Foundation models are trained on very large datasets and typically have billions of parameters. The computational requirements for training are massive, needing thousands of GPUs/TPUs and costing substantial energy resources. Moreover, extensive computing for training and inference has huge environmental impacts without conscientious mitigation efforts, especially carbon emissions. Costs and computing requirements concentrate progress in few organizations, risking monopoly power over impactful technologies.
To enhance our comprehension of the intricate diversity within geoscience data, a systematic classification encompassing space, air, ground, and ocean realms is imperative. This classification scheme offers a refined perspective on the multifaceted nature of these data types. The ensuing enumeration delineates prevalent data categories prevalent within the field: (i) Satellite and remote sensing data: Geoscientists extensively leverage satellite imagery and remote sensing data, which encompass optical, infrared, radar, and lidar datasets [88]. These sources provide invaluable insights into land cover, vegetation distribution, topographical variations, geological formations, and dynamic environmental changes. (ii) Geological and geophysical data: Geoscientists gather and analyze a plethora of geological and geophysical data, including seismic readings, gravity, and magnetic measurements, well logs, core samples, and geological maps [89]. These datasets are fundamental in decoding subsurface structures, discerning rock properties, and comprehending geological processes. (iii) Climate and weather data: Comprising parameters like temperature, precipitation patterns, humidity levels, solar radiation, wind dynamics, and atmospheric pressure, climate and weather data [90], which play a pivotal role in unraveling climate change phenomena, enhancing weather forecasting accuracy, and illuminating intricate interactions within the atmosphere, land, cryosphere and oceans. (iv) Geospatial data: Geospatial data integration, facilitated through Geographic Information Systems (GIS), harmonizes diverse Earth science datasets such as georeferenced maps, spatial databases, and geographic coordinates [91]. This integration empowers spatial analysis, visual representation, and predictive modeling. (v) Environmental monitoring data: This category encapsulates essential parameters like air and water quality, soil composition, and pollutant concentrations [92]. Environmental monitoring data is instrumental in scrutinizing shifts in environmental conditions, conducting comprehensive pollution assessments, and devising sustainable resource management strategies. (vi) Time-series data: Time-series observations, spanning climate records, seismic datasets, flow measurements, and glacier motion data, provide a chronicle of temporal changes [42]. Time-series analysis offers a means to identify patterns, track trends, and detect anomalies within these dynamic datasets. (vii) Text and scientific literature data: Geoscientists extensively tap into textual resources extracted from scientific literature, research publications, field reports, and scholarly notes [1]. These textual reservoirs not only furnish valuable insights but also serve as crucial references for guiding and informing ongoing scientific investigations.
IV-C Interpretability
Foundation models like GPT-4 are difficult to interpret because of their scale and complexity. With billions of parameters, it is challenging to explain why they generate a particular output. The inner workings seem opaque compared to smaller, more specialized AI systems. On the other hand, foundation models have some degree of interpretability. Techniques like attention mechanisms provide some visibility into what input the model is focusing on. There are also methods to analyze feature importance and conduct ablation studies to understand how different components of the model contribute to its outputs. The natural language capabilities of models like GPT-3 also lend themselves to certain interrogation techniques. Overall, there are challenges in interpreting large foundation models, but there are also opportunities. Interpretability is an active area of research, and new techniques are being developed. There is always a spectrum of interpretability, where no model is perfectly transparent or perfectly opaque. Researchers continue to explore methods to open the “black box” of foundation models while also recognizing their limitations. There are reasonable arguments on multiple sides of this complex issue.
IV-D Domain Knowledge
The complexity of geoscience domain knowledge arises from several intrinsic challenges. Firstly, it spans a broad spectrum of disciplines, including geology, geophysics, meteorology climate science, oceanography, and seismology, among others. Given that much of the Earth’s interior and history remains inaccessible for direct observation, integrating knowledge across these diverse fields is inherently complex. Secondly, geoscience operates from sub-daily to geological timescales, encompassing approximately 4.5 billion years of Earth system evolution. This extensive temporal scope means that distilling even fundamental knowledge becomes an immensely challenging task. Thirdly, the Earth system is characterized by intricate interconnections between the geosphere, hydrosphere, cryosphere, atmosphere, and biosphere. Understanding the dynamics of these interconnected mechanisms requires a holistic, interdisciplinary perspective. Furthermore, geoscience phenomena occur across a vast range of scales, from microscopic to planetary, making direct observation and sampling challenging. Lastly, the field of geoscience is continuously evolving, with new discoveries frequently expanding, revising, and refining our understanding of the Earth’s processes. Consequently, summaries and overviews can quickly become outdated as knowledge advances. This dynamic nature of geoscience knowledge further adds to the complexity of defining a comprehensive domain understanding.
IV-E Transparency and Trust
Previous research and reactions within scientific communities have revealed that the implementation of AI in geosciences raises concerns among geoscientists. Specifically, GFMs are often regarded as black boxes, presenting challenges in comprehending the rationale behind their predictions or explanations [93]. Geoscientists who emphasize transparency in their research may have reservations about using GFMs. Moreover, geoscientists working in data-scarce fields or those facing limited access to pertinent datasets may encounter challenges in the applicability and effectiveness of GFMs. When the available dataset inadequately represents specific scenarios or harbors harmful correlations, the risk of bias in test results during the training of GFMs becomes more pronounced. Ensuring the absence of undesirable biases in these unprecedentedly large and complex training datasets poses a formidable task, with the potential for social biases to amplify alongside the scale of the model. Additionally, geoscientists may be concerned about over-reliance on AI, which could result in a detrimental erosion of human expertise and critical thinking within the decision-making process.
GFMs necessitate rigorous verification across diverse scenarios prior to their promotion and application, guaranteeing their robust performance in critical and specialized contexts. Moreover, ongoing auditing and governance are imperative even after the models are deployed, considering the emergence of new challenges as they encounter novel tasks and settings. Notably, incentivizing the AI community through prize-winning competitions can serve as a catalyst for meticulous scrutiny of GFMs. For instance, participants could be rewarded for identifying indicators that engender harmful content or expose potential failure modes. Swiftly identifying and rectifying biases must remain a paramount objective for developers, suppliers, and regulators alike.
V Conclusion
GFMs hold immense promise for a significant advancement in modeling and prediction of Earth’s systems dynamics, thus enhancing our understanding of these intricate processes. Operating as a groundbreaking paradigm, GFMs inherit the idea of data-centric AI by leveraging large datasets and computational power to uncover concealed insights. Through flexible task specification, multi-modal inputs and outputs, and multi-modal knowledge, these GFMs models offer comprehensive insights that are infeasible through individual data sources or methodologies. Most importantly, the scalability and generalizability of GFMs make them uniquely capable of tackling a wide range of prediction, simulation, and decision-making challenges related to the Earth system dynamics, encompassing the highly nonlinear components and their complex multi-scale interactions. Going forward, fostering collaboration across various Earth science disciplines, in conjunction with computer scientists and applied mathematicians, will be pivotal in ushering in the next generation of GFMs. With their unique ability to explore multiple scenarios and quantify uncertainties, these models promise to provide critical insights into pressing issues such as climate change, natural hazard assessment, and sustainable resource management. Recent advancements in GFMs have demonstrated their potential to revolutionize the field of Earth system modeling, especially in the context of climate and weather forecasting, remote sensing, and geoscience. Despite prevailing challenges in validation and verification, scalability, interpretability, domain knowledge, and addressing social biases, GFMs have become an indispensable part of analyzing the past, monitoring the present, and planning for the future of the Earth.

### v4

@@ IV-C3 Applications and Disscussion
Datasets[253, 264, 265] Image-text datasets - - Remote sensing RSICap and RSIEval datasets for VLMs
RemoteCLIP [256] Unified image-caption data Hybrid Fine-tuning Various downstream tasks First general-purpose VLM for RS
CLIP-RS [257] RS images and text Contrastive Fine-tuning Cross-modal RS image retrieval CLIP
H2RSVLM [258] RS images with captions - - Large scale captions for RS images 1.4 million image-caption pairs
Al Rahhal et al. [259] RS images with captions - Prompt tuning Zero-shot RS classification CLIP/Open-CLIP
VRSBench [255] Images, captions, question-answer pairs - - Remote sensing images A new benchmark
He et al. [260] Multimodal RS datasets - - Multimodal RS data classification Foundation model adaptation (FMA) framework
Prompt-RSVQA [266] Visual images with questions and answers Hybrid Prompt tuning VQA in remote sensing -
RS-CapRet [267] Remote sensing images Contrastive Fine-tuning Captioning and retrieving remote sensing images -
Text2seg [268] Remote sensing images - Prompt Tuning Remote sensing image segmentation -
V Challenges and Future Trends
In this section, we discuss the challenges and future trends in GFMs, including data integration, model complexity, uncertainty quantification, interdisciplinary collaboration, and privacy, trust and security concerns. The challenges and future trends in GFMs are summarized in Fig. 16.
Figure 16: Challenges and Future Trends in GFMs.
V-A Data Integration
Effective data integration is pivotal for the development of reliable GFMs. Addressing challenges associated with data diversity, quality, and volume is essential for GFMs to provide accurate and actionable insights into Earth’s intricate systems. In geoscience, the integration of diverse datasets into a coherent, unified model that accurately represents these systems poses a significant challenge. This process is further complicated by the heterogeneity of data sources, variable data quality, and the substantial volumes of data produced by modern sensing technologies.
V-A1 Variety of Data Sources
Geoscientific data originates from a variety of sources, including satellite observations, geological surveys, atmospheric sensors, and oceanographic instruments [269, 270, 271, 272, 273, 65]. These sources produce data in diverse formats, scales, and resolutions, complicating their integration into a unified model. For instance, satellite data may offer extensive spatial coverage but lack the temporal resolution needed for real-time monitoring. In contrast, ground-based sensors can provide high-frequency data that is, however, geographically restricted [274].
V-A2 Data Quality and Compatibility
The quality of geoscience data varies significantly depending on the source and collection method [22]. Common issues such as data inaccuracies, missing values, and measurement errors can introduce significant biases into model outputs if not adequately addressed. Moreover, ensuring data compatibility across different scales and formats necessitates sophisticated preprocessing techniques, such as normalization, error correction, and interpolation, to guarantee meaningful contributions from all data inputs to the models.
V-A3 Volume and Scalability
Foundation models, trained on extensive datasets, typically possess billions of parameters [30]. These models demand substantial computational resources for training, requiring thousands of GPUs or TPUs and consuming considerable energy. Additionally, the rapid increase in data produced by digital technologies poses a significant challenge. Managing and processing these large datasets necessitates advanced computational resources and sophisticated data management systems. Scalability is essential, as models must efficiently process increasing volumes of data without compromising processing speed or accuracy [275].
V-A4 Integration Techniques
Addressing these challenges often entails employing advanced data integration techniques, including sophisticated machine learning algorithms that automate data cleaning, integration, and analysis [276]. Geographic Information Systems (GIS) and other spatial analysis tools are crucial in aligning disparate data types across spatial dimensions [270]. Additionally, data assimilation techniques, traditionally used in meteorology and oceanography to systematically combine model predictions with observational data, are being adapted for broader geoscience applications to improve model accuracy and reliability.
V-B Model Complexity
Addressing the complexities inherent in GFMs is essential to enhancing their functional capabilities. Critical areas for ongoing research and development encompass streamlining models to retain essential details effectively, advancing computational techniques, and enhancing stakeholder communication. GFMs are designed to simulate the intricate interactions within Earth’s systems across various temporal and spatial scales, a complexity that poses substantial challenges in model development, computational requirements, and the interpretation of outputs.
V-B1 Multi-Scale Modeling
GFMs require the capacity to operate effectively across a range of scales. Specifically, global climate models must integrate data and processes at the planetary scale while concurrently delivering accurate predictions at regional and local levels [277]. This multi-scale requirement complicates the modeling process, as each scale may necessitate unique modeling approaches and distinct data resolutions to accurately reflect the varying dynamics and interactions [278].
V-B2 Simulating Complex Processes
Earth systems are composed of a wide range of dynamic processes, including atmospheric dynamics, hydrological cycles, geological transformations, and ecological interactions [3, 279]. Each process exhibits significant inherent complexity, yet primary challenges stem from accurately modeling the interactions among these diverse elements. For instance, effectively predicting the impacts of climate change on river basins requires a deep understanding not only of atmospheric physics and hydrology but also of relevant human activities [280]. This necessitates an integrated approach that encompasses multiple scientific disciplines to ensure that models can accurately reflect the multifaceted nature of earth systems.
V-B3 Computational Demands
The complexity inherent in GFMs requires extensive computational resources. These models typically incorporate advanced numerical methods, utilize high-resolution data, and necessitate prolonged simulation periods, all demanding substantial computational power. As a result, high-performance computing (HPC) systems are often indispensable, leading to considerable expenses and requiring specialized technical expertise [281]. This reliance on HPC underscores the critical need for optimization and efficiency improvements in model algorithms to manage costs and accessibility issues effectively.
V-B4 Validation and Verification
Unlike conventional geoscience AI models, which are designed for specific tasks and validated within predefined use case parameters, GFMs are engineered to perform novel tasks as specified by end-users without additional training [30]. This capability considerably complicates predicting potential failure modes in such versatile systems. Effective testing of GFMs requires accumulating illustrative cases over extended periods to statistically capture rare events [15, 17]. Furthermore, as these models become more complex, the challenge of ensuring accurate predictions intensifies. Validating such models often necessitates extensive field data, which can be difficult and costly to obtain. Additionally, with increasing complexity, it becomes progressively challenging to identify the components responsible for inaccuracies in the outputs.
V-B5 Model Interpretability
Interpretability remains an active field of research, continuously evolving with the development of new techniques. The degree of interpretability varies, as no model is entirely transparent or utterly opaque. Researchers are consistently working on methods to demystify the ”black box” nature of foundation models while recognizing their inherent limitations [29]. Foundation models, such as ChatGPT-4 and Claude-3.5, present interpretative challenges due to their substantial scale and complexity [16, 40]. Possessing billions of parameters, these models make it challenging to elucidate the reasons behind their specific outputs. Consequently, the complexity of GFMs can hinder stakeholders, including policymakers and the general public, from fully comprehending and utilizing the insights these models provide. This challenge may prevent the effective integration of scientific discoveries into decision-making processes crucial for environmental management and policy formulation.
V-C Uncertainty Quantification
Effective uncertainty quantification is crucial for enhancing the credibility and utility of GFMs. Ongoing advancements in statistical modeling, computational techniques, and stakeholder communication are vital for managing the inherent uncertainties in modeling Earth systems [282, 283]. Accurate quantification of uncertainty in GFMs is essential for delivering reliable predictions and informing risk management and decision-making processes. However, the inherent unpredictability of natural systems, combined with limitations in data and modeling techniques, presents significant challenges [3]. Ongoing research in uncertainty quantification of GFMs is dedicated to devising more sophisticated methods that can incorporate various types of uncertainties, thereby offering more accurate and comprehensive risk assessments. Enhancements in computational power and data analytics facilitate more intricate uncertainty analyses, which are essential for augmenting the reliability and applicability of GFMs.
V-C1 Sources of Uncertainty
Uncertainties in GFMs stem from multiple sources, such as incomplete or inaccurate data, simplified assumptions in the modeling, and the inherent randomness of natural processes. Specifically, the prediction of earthquake occurrences illustrates these complexities well. Uncertainties here often arise from an incomplete understanding of subsurface geological conditions combined with the inherently chaotic nature of seismic processes [284, 285]. This multifaceted origin of uncertainties necessitates sophisticated modeling techniques to enhance accuracy and reliability in predictions.
V-C2 Statistical Methods for Uncertainty Quantification
Geoscientists utilize a variety of statistical methods to tackle uncertainties in their models. Probabilistic models, notably Bayesian networks, are employed to embed uncertainty directly within the model predictions [286]. These models yield a spectrum of possible outcomes rather than a singular deterministic result, thus providing a more nuanced understanding of potential scenarios. Additionally, Monte Carlo simulations represent another widely used approach [287]. This technique involves repeatedly running the model with varied inputs to produce a distribution of potential outcomes, thereby capturing the variability and uncertainty inherent in natural systems.
V-C3 Ensemble Modeling
Ensemble modeling has become a fundamental technique in fields such as climate science and hydrology [288]. This method involves running multiple distinct models or a single model under varying parameter settings to explore a wider array of scenarios. Such a comprehensive approach enables scientists to better capture the uncertainties inherent in system responses. Additionally, ensemble modeling facilitates the identification of predictions that are consistent across different models and those that are particularly sensitive to specific assumptions. This dual capability enhances the reliability of model outputs and aids in refining model assumptions based on comparative robustness.
V-C4 Communicating Uncertainty
A significant challenge in the field of geoscience involves effectively communicating the uncertainties associated with model predictions to non-specialist stakeholders. It is imperative for decision-makers to comprehend the level of confidence in model outputs to facilitate informed decision-making. This necessitates not only the transparent reporting of uncertainties but also proactive education and engagement with the users of model data. Such interactions help ensure that stakeholders fully understand the implications of uncertainty within the context of their specific decision-making processes.
V-D Interdisciplinary Collaboration
The development of comprehensive GFMs requires collaboration across multiple scientific disciplines, including climatology, oceanography, ecology, and geology. Effective interdisciplinary collaboration, although challenging due to the diversity of the disciplines involved, is crucial for advancing GFMs. It enhances the models’ comprehensiveness and accuracy by integrating a broad spectrum of scientific insights and methodologies. Key strategies to overcome barriers to effective interdisciplinary work include improving communication, integrating diverse models, and developing supportive collaborative platforms.
V-D1 Differences in Methodologies
Each scientific discipline contributes unique methodologies, tools, and perspectives to ensemble modeling, posing challenges in creating a cohesive modeling effort. For example, integrating data and models from atmospheric scientists with those from hydrologists requires the alignment of different scales, resolutions, and data types commonly employed in each field [289]. This integration is crucial for constructing comprehensive models that accurately represent complex interactions across disciplines, thereby enhancing the reliability and utility of the predictions.
V-D2 Communication Barriers
Effective communication across different scientific disciplines is frequently obstructed using discipline-specific jargon and conceptual differences. Terms and concepts often carry distinct meanings in various fields, which can lead to misunderstandings and misinterpretations. Establishing a common language and understanding is crucial for successfully integrating interdisciplinary insights into GFMs. This foundational communication strategy not only enhances collaboration but also significantly improves the accuracy and effectiveness of the models developed.
V-D3 Collaborative Platforms
Establishing effective platforms for collaboration is essential for advancing interdisciplinary research. This requires implementing both technological and organizational solutions. Technologically, shared databases and integrated modeling tools are vital for seamless data exchange and analysis [290]. Organizationally, frameworks that promote regular interaction and cooperation among researchers from diverse fields are equally important. Such structures should be designed to support continuous knowledge sharing and collective problem-solving, thereby enhancing the productivity and impact of collaborative efforts.
V-D4 Funding and Institutional Supports
Securing funding for interdisciplinary projects presents significant challenges, as traditional funding sources are often aligned with specific disciplines rather than cross-disciplinary initiatives. Furthermore, robust institutional support is essential to cultivate an environment that not only encourages but also values interdisciplinary collaboration [1]. Institutions must actively develop policies and provide resources that facilitate the integration of diverse scientific perspectives, thereby enhancing the scope and impact of research outcomes.
V-E Privacy, Trust and Security Concerns
As GFMs increasingly rely on large datasets containing potentially sensitive information, addressing concerns related to privacy, trust, and security becomes crucial. These issues are of particular importance when GFMs are deployed in contexts that could significantly impact public policies or individual lives. To advance and secure broader acceptance of GFMs, it is essential to effectively manage these concerns. The geoscience community must adopt stringent data management practices, enhance transparency, and implement robust security measures. Such efforts will not only improve the trustworthiness and ethical standing of GFMs but also facilitate their responsible and beneficial application in society.
V-E1 Data Privacy
GFMs frequently utilize data that may include geographically and personally identifiable information, such as land use patterns, water usage, or individual movements derived from mobile data. It is imperative to ensure that such data is handled in a manner that respects privacy rights. This commitment involves strict adherence to data protection regulations, such as the General Data Protection Regulation (GDPR) in Europe [291]. Additionally, the implementation of robust data anonymization and encryption techniques is essential to safeguard privacy. These measures are crucial for maintaining the integrity and trustworthiness of GFMs, thereby enabling their ethical use in research and policy-making.
V-E2 Trust in Models
The reliability of GFMs is crucial for establishing trust among diverse stakeholders, including the public, policymakers, and scientific communities. Trust in these models largely depends on the transparency of the modeling processes, the precision of the outputs, and clear communication regarding the models’ capabilities and limitations [292]. Establishing such trust is vital because misplaced confidence in flawed models can result in poor decision-making, leading to potentially severe repercussions. Therefore, continuous efforts to enhance model accuracy, process transparency, and stakeholder communication are essential to foster a robust foundation of trust.
V-E3 Security Concerns
The growing interconnectivity of modeling tools and data repositories heightens security risks within the geoscience field. Potential cyberattacks could result in data breaches, manipulation of model outputs, or disruptions to essential modeling infrastructure. To secure GFMs effectively, it is crucial to implement comprehensive cybersecurity measures. This security framework should include regular security audits and the development of robust contingency plans to ensure data integrity and continuous model operation. These proactive measures are fundamental to protecting the integrity of geoscience data and the reliability of model outputs.
V-E4 Ethical Use of Models
The ethical implications of model usage and accessibility extend beyond privacy and security concerns. It is crucial that models, especially those influencing significant economic or policy decisions, are developed and deployed responsibly. This responsibility entails careful attention to potential biases in the data or the models themselves, which could result in unfair or harmful outcomes. Ensuring ethical practices involves maintaining transparency in model development processes, rigorously testing for biases, and making these models accessible to a broad range of stakeholders to avoid privileging or disadvantaging any particular group.
V-E5 Legal and Regulatory Compliance
As the application of geoscience models becomes more widespread, adhering to a complex array of international, national, and local regulations poses a significant challenge. These regulations often cover aspects such as data usage, environmental assessments, and the public dissemination of predictive results. For geoscience modeling efforts to maintain their legitimacy and effectiveness, it is crucial to stay well-informed of relevant laws and ensure strict compliance. This includes regularly updating compliance protocols and training for model developers and users to navigate the evolving legal landscape effectively.
VI Conclusion
Geoscience Foundation Models (GFMs) represent a paradigm shift in Earth system modeling, leveraging data-centric AI to integrate massive cross-disciplinary datasets and address complexities that traditional models struggle to capture. This review has highlighted GFMs’ unique strengths in flexible task specification, diverse input-output capabilities, and multi-modal knowledge representation, while also presenting key construction techniques focusing on transformers, pre-training, and adaptation. The exploration of recent advances in large language, vision, and vision-language models, particularly in remote sensing applications, demonstrates GFMs’ broad potential in geoscience. However, significant challenges remain in data integration, model complexity, uncertainty quantification, interdisciplinary collaboration, and addressing privacy, trust, and security concerns. As we move forward, addressing these challenges will be crucial to fully realizing GFMs’ potential. The future of GFMs lies in their continued evolution towards more integrated, data-driven modeling approaches, offering a more comprehensive understanding of Earth’s complex systems. This progression promises to provide critical insights into pressing global issues such as climate change, natural hazards, and sustainability, ultimately transforming our approach to Earth system science and informing decision-making in the face of global environmental challenges.